OpenStudy (anonymous):

solve the equation √2x+1=√x+4 a. 3 b. 1/3 c. -3 d. 1

OpenStudy (anonymous):

@dan815 @nincompoop @Michele_Laino

OpenStudy (anonymous):

\[\sqrt{2x+1}=\sqrt{x+4}\] like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what's stopping you from just squaring both sides then?

OpenStudy (anonymous):

Just one thing, if you square away a root in an equation, always make sure you check that any numbers that appear to be a valid answer do not make the outcome of the equation that would have been under a root negative. Because \(\sqrt{-8}\) for example is not a real number (unless you're talking imaginary numbers)